Before the government decreased the tax rate, how much of government
spending was required to bring the economy to full employment?
1
Expert's answer
2020-12-17T04:49:34-0500

This task is part of the following problem:

The following information is provided about an open economy with a government. Use the information to answer the questions that follow:

C = 450 + 0.4Y

I = 350

G = 150

X = 70

Z = 35 + 0.1Y

T = 0.15Y

Yf = 1550

C= Consumption spending bu households

G = Government spending

X= Receipts for exports

Z= Payment for imports

Before the government decreased the tax rate, how much government spending was required to bring the economy to full employment?

For this, You need the initial aggregate spending with taxes.

"AE = 450+0.4(Y-0.15Y) + 350 + 150 + 70 -35 - 0.1Y \\

= 985 + 0.4 \times 0.85Y -0.1Y \\

= 985 + 0.34Y-0.1Y \\

= 985 + 0.24Y"

So equilibrium

"Y = 985 +0.24Y \\

0.76Y = 985"

Y** = 1296.05

The gap between initial output and full employment output is Yf-Y** = 253.95

Multiplier for the economy is

"\frac{1}{leakages} = \frac{1}{0.76} = 1.31"

So a unit change in government spending will raise the output by 1.31 times. So to close the output hole of 253.95 and rise in government spending is needed by the measure of

"\frac{253.95}{1.31} = 193.85"

So absolute government spending should be 343.85.